CONSIDER THE FOLLOWING SHORT RUN PRODUCTION FUNCTION Q=6L2-0.4L3
a. find the value of L that maximizes out put
b find the value of L that maximizes MARGINAL PRODUCT
c. find the value of l that maximizes average product
1
Expert's answer
2020-06-08T11:10:23-0400

a) To maximize the output we have to solve the equation:

"Q'=(6L\u00b2-0.4L\u00b3)'= 0"

"12L-1.2L\u00b2=0"

"L(12-1.2L)=0"

"L1=0"

"L2=10"

b)"MP=12L-1.2L\u00b2." To maximize we have to calculate: MP'=0, so

"12-2.4L=0"

"L=5."

c)"AP=6L-0.4L\u00b2" To maximize we have to calculate AP'=0, so

"6-0.8L=0"

"L=7.5"